Proxy Computing Co. has the following costs: TC = 256 + 128Q + 8Q2

(i) Identify the fixed, variable and marginal costs. Does the cost structure represent a short-run or long-run cost structure? Why?

(ii) The market demand curve is: Q = 1000 – 2P and the supply curve is: Q = 3P. Proxy Computing is a small firm operating in this perfectly competitive market. Compute Proxy Computing’s profit-maximising quantity.
Answer #1

(i)

FC = 256

TVC = 128Q + 8Q2

MC = dTC/dQ = 128 + 16Q

Since FC > 0, this is short run, since in long run FC = 0.

(ii)

Equating market demand and supply,

1000 - 2P = 3P

5P = 1000

P = 200

Q = 3 x 200 = 600

The firm equates market price with its MC:

128 + 16Q = 200

16Q = 72

Q = 4.5 (firm output)